# Get rich quick: exact solutions reveal how unbalanced initializations promote rapid feature learning

### v1

@@ Abstract
While the impressive performance of modern neural networks is often attributed to their capacity to efficiently extract task-relevant features from data, the mechanisms underlying this rich feature learning regime remain elusive, with much of our theoretical understanding stemming from the opposing lazy regime. In this work, we derive exact solutions to a minimal model that transitions between lazy and rich learning, precisely elucidating how unbalanced layer-specific initialization variances and learning rates determine the degree of feature learning. Our analysis reveals that they conspire to influence the learning regime through a set of conserved quantities that constrain and modify the geometry of learning trajectories in parameter and function space. We extend our analysis to more complex linear models with multiple neurons, outputs, and layers and to shallow nonlinear networks with piecewise linear activation functions. In linear networks, rapid feature learning only occurs with balanced initializations, where all layers learn at similar speeds. While in nonlinear networks, unbalanced initializations that promote faster learning in earlier layers can accelerate rich learning. Through a series of experiments, we provide evidence that this unbalanced rich regime drives feature learning in deep finite-width networks, promotes interpretability of early layers in CNNs, reduces the sample complexity of learning hierarchical data, and decreases the time to grokking in modular arithmetic. Our theory motivates further exploration of unbalanced initializations to enhance efficient feature learning.
@@ 1 Introduction
Lazy regime. Various investigations have revealed a notable phenomenon in overparameterized neural networks, where throughout training the networks remain close to their linearization [1, 2, 3, 4, 5]. Seminal work by Jacot et al. [6], demonstrated that in the infinite-width limit, the Neural Tangent Kernel (NTK), which describes the evolution of the neural network through training, converges to a deterministic limit. Consequently, the network learns a solution akin to kernel regression with the NTK matrix. Termed the lazy or kernel regime, this domain has been characterized by a deterministic NTK [6, 7], convex dynamics with minimal movement in parameter space [8], static hidden representations, exponential learning curves, and implicit biases aligned with a reproducing kernel Hilbert space (RKHS) norm [9]. However, Chizat et al. [8] challenged this understanding, asserting that the lazy regime isn’t a product of the infinite-width architecture, but is contingent on the scale of the network at initialization. They demonstrated that given any finite-width model $f(x;\theta)$ whose output is zero at initialization, a scaled version of the model $\tau f(x;\theta)$ will enter the lazy regime as the scale $\tau$ diverges. However, they also noted that these scaled models often perform worse in test error. While the lazy regime offers insights into the network’s convergence to a global minimum, it does not fully capture the generalization capabilities of neural networks trained with standard initializations. It is thus widely believed that a different regime, driven by small or vanishing initializations, underlies the many successes of neural networks.
@@
Rich regime. In contrast to the lazy regime, the rich or feature-learning or active regime is distinguished by a learned NTK that evolves through training, non-convex dynamics traversing between saddle points [10, 11, 12], sigmoidal learning curves, and simplicity biases such as low-rankness [13] or sparsity [14]. Yet, the exact characterization of rich learning and the features it learns frequently depends on the specific problem at hand, with its definition commonly simplified as what it is not: lazy. Recent analyses have shown that beyond scale, other aspects of the initialization can substantially impact the extent of feature learning, such as the effective rank [15], layer-specific initialization variances [16, 17, 18], and large learning rates [19, 20, 21, 22]. Azulay et al. [9] demonstrated that in two-layer linear networks, the relative difference in weight magnitudes between the first and second layer, termed the initialization geometry in our work, can impact feature learning, with balanced initializations yielding rich learning dynamics, while unbalanced ones tend to induce lazy dynamics. However, as shown in Fig. 1, for nonlinear networks unbalanced initializations can induce both rich and lazy dynamics, creating a complex phase portrait of learning regimes influenced by both scale and geometry. Building on these observations, our study aims to precisely understand how layer-specific initialization variances and learning rates determine the transition between lazy and rich learning in finite-width networks. Moreover, we endeavor to gain insights into the inductive biases of both regimes, and the transition between them, during training and at interpolation, with the ultimate goal of elucidating how the rich regime acquires features that facilitate generalization.
(a) Scale and geometry influence the learning regime
@@
Figure 1: Unbalanced initializations lead to rapid rich learning and generalization. We follow the experimental setup used in Fig. 1 of Chizat et al. [8] – a wide two-layer student ReLU network $f(x;\theta)=\sum_{i=1}^{h}a_{i}\max\{0,w_{i}^{\intercal}x\}$ trained on a dataset generated from a narrow two-layer teacher ReLU network. The student parameters are initialized as $w_{i}\sim\text{Unif}(\mathbb{S}^{d-1}(\frac{\tau}{\alpha}))$ and $a_{i}=\pm\alpha\tau$ such that $\tau>0$ controls the scale of the function at initialization, while $\alpha>0$ controls the geometry, the relative magnitude of the first and second layer parameters. The conserved quantity is $\delta=\tau^{2}(\alpha^{2}-\alpha^{-2})$. (a) Shows the training trajectories of $|a_{i}|w_{i}$ (color denotes $\mathrm{sgn}(a_{i})$) when $d=2$ for four different settings of $\tau,\delta$. The left plot confirms that small scale leads to rich and large scale to lazy. The right plot shows that even at small scale, the initialization geometry can move the network between rich and lazy as well. Here an upstream initialization $\delta>0$ shows striking alignment to the teacher (dotted lines), while a downstream initialization $\delta<0$ shows no alignment. (b) Shows the test loss and kernel distance from initialization computed through training over a sweep of $\tau$ and $\delta$ when $d=100$. Lazy learning happens when $\tau$ is large, rich learning happens when $\tau$ is small, and rapid rich learning happens when both $\tau$ is small and $\delta$ is large – an upstream initialization. This initialization also leads to the smallest test loss. See Fig. 10 in Section D.1 for supporting figures.
@@
Our contributions. Our work begins with an exploration of the two-layer single-neuron linear network proposed by Azulay et al. [9] as a minimal model displaying both lazy and rich learning. By employing a combination of hyperbolic and spherical coordinate transformations, we derive exact solutions for the gradient flow dynamics with layer-specific learning rates under all initializations. Alongside recent work by Xu and Ziyin [23] [^1] Xu and Ziyin [23] presented exact NTK dynamics for a linear model trained with one-dimensional data., our analysis stands out as one of the few analytically tractable models for the transition between lazy and rich learning in a finite-width network, marking a notable contribution to the field. Our analysis reveals that the layer-specific initialization variances and learning rates, which we collectively refer to as the initialization geometry, conspire to influence the learning regime through a simple set of conserved quantities that constrain the geometry of learning trajectories. Additionally, it reveals that a crucial aspect of the initialization geometry overlooked in prior analysis is the directionality. While a balanced initialization results in all layers learning at similar rates, an unbalanced initialization can cause faster learning in either earlier layers, referred to as an upstream initialization, or later layers, referred to as a downstream initialization. Due to the depth-dependent expressivity of layers in a network, upstream and downstream initializations often exhibit fundamentally distinct learning trajectories. We extend our analysis of the initialization geometry developed in the single-neuron model to more complex linear models with multiple neurons, outputs, and layers and to two-layer nonlinear networks with piecewise linear activation functions. We find that in linear networks, rapid rich learning can only occur with balanced initializations, while in nonlinear networks, upstream initializations can actually accelerate rich learning. Finally, through a series of experiments, we provide evidence that upstream initializations drive feature learning in deep finite-width networks, promote interpretability of early layers in CNNs, reduce the sample complexity of learning hierarchical data, and decrease the time to grokking in modular arithmetic.
@@
Notation. In this work, we consider a feedforward network $f(x;\theta):\mathbb{R}^{d}\to\mathbb{R}$ parameterized by $\theta\in\mathbb{R}^{p}$. The network is trained by gradient flow $\dot{\theta}_{i}=-\eta_{\theta_{i}}\nabla_{\theta_{i}}\mathcal{L}(\theta)$, with an initialization $\theta_{0}$ and layer-specific learning rate $\eta_{\theta_{i}}>0\in\mathbb{R}^{p}$, to minimize the mean squared error $\mathcal{L}(\theta)=\frac{1}{2}\sum_{i=1}^{n}(f(x_{i};\theta)-y_{i})^{2}$ computed over a dataset $\{(x_{1},y_{1}),\dots,(x_{n},y_{n})\}$ of size $n$. We denote the training data matrix as $X\in\mathbb{R}^{n\times d}$ with rows $x_{i}\in\mathbb{R}^{d}$ and the label vector as $y\in\mathbb{R}^{n}$. The network’s output $f(x;\theta)$ evolves according to the differential equation, $\partial_{t}f(x;\theta)=-\sum_{i=1}^{n}\Theta(x,x_{i};\theta)(f(x_{i};\theta)-y_{i})$, where $\Theta(x,x^{\prime};\theta):\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ is the Neural Tangent Kernel (NTK), defined as $\Theta(x,x^{\prime};\theta)=\sum_{i=1}^{p}\eta_{\theta_{i}}\partial_{\theta_{i}}f(x;\theta)\partial_{\theta_{i}}f(x^{\prime};\theta)$. The NTK quantifies how one gradient step with data point $x^{\prime}$ affects the evolution of the networks’s output evaluated at another data point $x$. When $\eta_{\theta_{i}}$ is shared by all parameters, the NTK is the kernel associated with the feature map $\nabla_{\theta}f(x;\theta)\in\mathbb{R}^{p}$. We also define the NTK matrix $K\in\mathbb{R}^{n\times n}$, which is computed across the training data such that $K_{ij}=\Theta(x_{i},x_{j};\theta)$. The NTK matrix evolves from its initialization $K_{0}$ to convergence $K_{\infty}$ through training. Lazy and rich learning exist on a spectrum, with the extent of this evolution serving as the distinguishing factor. Various studies have proposed different metrics to track the evolution of the NTK matrix [24, 25, 26, 27]. In this work, we use kernel distance [25], defined as $S(t_{1},t_{2})=1-\langle K_{t_{1}},K_{t_{2}}\rangle/\left(\|K_{t_{1}}\|_{F}\|K_{t_{2}}\|_{F}\right)$, which is a scale invariant measure of similarity between the NTK at two points in time. In the lazy regime $S(0,t)\approx 0$, while in the rich regime $0\ll S(0,t)\leq 1$.
@@ 2 Related Work
Infinite-width networks. Many recent advancements in understanding the rich regime have come from studying how the initialization variance and layer-wise learning rates should scale in the infinite-width limit to ensure constant movement in the activations, gradients, and outputs. In this limit, analyzing dynamics becomes simpler in several respects: random variables concentrate, nonlinearities act linearly, and quantities will either vanish to zero, remain constant, or diverge to infinity [17]. A set of works used tools from statistical mechanics to provide analytic solutions for the rich population dynamics of two-layer nonlinear neural networks initialized according to the mean field parameterization [56, 57, 58, 59]. These ideas were extended to deeper networks through a tensor program framework, leading to the derivation of maximal update parametrization ($\mu\mathrm{P}$) [16, 18]. The $\mu\mathrm{P}$ parameterization has also been derived through a self-consistent dynamical mean field theory [60] and a spectral scaling analysis [61]. In this study, we focus on finite-width neural networks.
@@ 3 A Minimal Model of Lazy and Rich Learning with Exact Solutions
Deriving exact solutions in parameter space. We assume whitened input $X^{\intercal}X=\mathbf{I}_{d}$ such that the ordinary least squares solution is $\beta_{*}=X^{\intercal}y$, and the gradient flow dynamics simplify to $\dot{a}=\eta_{a}\left(w^{\intercal}\beta_{*}-a\|w\|^{2}\right)$, $\dot{w}=\eta_{w}\left(a\beta_{*}-a^{2}w\right)$. We note that $w(t)\in\mathrm{span}(\{w_{0},\beta_{*}\})$, and through training, $w$ aligns in direction to ±$\pm\beta_{*}$ depending on the basin of attraction [^3] The basin is given by the sign of $a_{0}$ for $\delta\geq 0$ or $w_{0}^{\intercal}\beta_{*}+\frac{a_{0}}{2}(\delta+\sqrt{\delta^{2}+4\|\beta_{*}\|^{2}})$ for $\delta<0$. See A.2.5. the parameters are initialized in. Therefore, we can monitor the dynamics by tracking the hyperbolic geometry between $a$ and $\|w(t)\|$ and the spherical angle between $w(t)$ and $\beta_{*}$. We study the variables $\mu=a\|w\|$, an invariant under the rescale symmetry, and $\phi=\frac{w^{\intercal}\beta_{*}}{\|w\|\|\beta_{*}\|}$, the cosine of the spherical angle. From these two scalar quantities $\mu(t),\phi(t)$ and the initialization $a_{0},w_{0}$, we can determine the trajectory $a(t)$ and $w(t)$ in parameter space. The dynamics for $\mu,\phi$ are given by the coupled nonlinear ODEs,
@@
Amazingly, this system can be solved exactly, as discussed in Section A.2, and shown in Fig. 3. Without delving into the specifics, we can develop an intuitive understanding of the solutions by examining the influence of the initialization geometry $\delta$.
@@
Identifying regimes of learning in function space. Here we take an alternative route towards understanding the influence of the initialization geometry by directly examining the dynamics in function space, an analysis strategy we will generalize to broader setups in Sections 4 and 5. The network’s function is determined by the product $\beta=aw$ and governed by the ODE,
@@
where $\rho=X\beta-y$ is the residual. These dynamics can be interpreted as preconditioned gradient flow on the loss in function space where the preconditioning matrix $M$ depends on time through its dependence on $a^{2}$ and $ww^{\intercal}$. Whenever $\|\beta\|\neq 0$, we can express $M$ in terms of $\beta$ and $\delta$ as
$M=\frac{\kappa+\delta}{2}I_{d}+\frac{\kappa-\delta}{2}\frac{\beta\beta^{\intercal}}{\|\beta\|^{2}},$ (4)
where $\kappa=\sqrt{\delta^{2}+4\eta_{a}\eta_{w}\|\beta\|^{2}}$ (see Section A.3 for a derivation). This establishes a self-consistent equation for the dynamics of $\beta$ regulated by $\delta$. Additionally, notice that $M$ characterizes the NTK matrix Eq. 1. Thus, understanding the evolution of $M$ along the trajectory $\beta_{0}$ to $\beta_{*}$ offers a method to discern between lazy and rich learning. Upstream. When $\delta\gg 0$, $M\approx\delta\mathbf{I}_{d}$, and the dynamics of $\beta$ converge to the trajectory of linear regression trained by gradient flow. Along this trajectory the NTK matrix remains constant, confirming the dynamics are lazy. Balanced. When $\delta=0$, $M=\sqrt{\eta_{a}\eta_{w}}\|\beta\|(\mathbf{I}_{d}+\tfrac{\beta\beta^{\intercal}}{\|\beta\|^{2}})$. Here the dynamics balance between following the lazy trajectory and attempting to fit the task by only changing in norm. As a result the NTK changes in both magnitude and direction through training, confirming the dynamics are rich. Downstream. When $\delta\ll 0$, $M\approx|\delta|\tfrac{\beta\beta^{\intercal}}{\|\beta\|^{2}}$, and $\beta$ follows a projected gradient descent trajectory, attempting to reach $\beta_{*}$ in the direction of $\beta_{0}$. Along this trajectory the NTK matrix doesn’t evolve. However, if $\beta_{0}$ is not aligned to $\beta_{*}$, then at some point the dynamics of $\beta$ will slowly align. In this second alignment phase the NTK matrix will change, confirming the dynamics are initially lazy followed by a delayed rich phase.

### v2

@@ Abstract
While the impressive performance of modern neural networks is often attributed to their capacity to efficiently extract task-relevant features from data, the mechanisms underlying this rich feature learning regime remain elusive, with much of our theoretical understanding stemming from the opposing lazy regime. In this work, we derive exact solutions to a minimal model that transitions between lazy and rich learning, precisely elucidating how unbalanced layer-specific initialization variances and learning rates determine the degree of feature learning. Our analysis reveals that they conspire to influence the learning regime through a set of conserved quantities that constrain and modify the geometry of learning trajectories in parameter and function space. We extend our analysis to more complex linear models with multiple neurons, outputs, and layers and to shallow nonlinear networks with piecewise linear activation functions. In linear networks, rapid feature learning only occurs from balanced initializations, where all layers learn at similar speeds. While in nonlinear networks, unbalanced initializations that promote faster learning in earlier layers can accelerate rich learning. Through a series of experiments, we provide evidence that this unbalanced rich regime drives feature learning in deep finite-width networks, promotes interpretability of early layers in CNNs, reduces the sample complexity of learning hierarchical data, and decreases the time to grokking in modular arithmetic. Our theory motivates further exploration of unbalanced initializations to enhance efficient feature learning.
@@ 1 Introduction
Lazy regime. Various investigations have revealed a notable phenomenon in overparameterized neural networks, where throughout training the networks remain close to their linearization [1, 2, 3, 4, 5]. Seminal work by Jacot et al. [6], demonstrated that in the infinite-width limit, the Neural Tangent Kernel (NTK), which describes the evolution of the neural network through training, converges to a deterministic limit. Consequently, the network learns a solution akin to kernel regression with the NTK matrix. Termed the lazy or kernel regime, this domain has been characterized by a deterministic NTK [6, 7], minimal movement in parameter space [8], static hidden representations, exponential learning curves, and implicit biases aligned with a reproducing kernel Hilbert space (RKHS) norm [9]. However, Chizat et al. [8] challenged this understanding, asserting that the lazy regime isn’t a product of the infinite-width architecture, but is contingent on the overall scale of the network at initialization. They demonstrated that given any finite-width model $f(x;\theta)$ whose output is zero at initialization, a scaled version of the model $\tau f(x;\theta)$ will enter the lazy regime as the scale $\tau$ diverges. However, they also noted that these scaled models often perform worse in test error. While the lazy regime offers insights into the network’s convergence to a global minimum, it does not fully capture the generalization capabilities of neural networks trained with standard initializations. It is thus widely believed that a different regime, driven by small or vanishing initializations, underlies the many successes of neural networks.
@@
Rich regime. In contrast to the lazy regime, the rich or feature-learning or active regime is distinguished by a learned NTK that evolves through training, non-convex dynamics traversing between saddle points [10, 11, 12], sigmoidal learning curves, and simplicity biases such as low-rankness [13] or sparsity [14]. Yet, the exact characterization of rich learning and the features it learns frequently depends on the specific problem at hand, with its definition commonly simplified as what it is not: lazy. Recent analyses have shown that beyond overall scale, other aspects of the initialization can substantially impact the extent of feature learning, such as the effective rank [15], layer-specific initialization variances [16, 17, 18], and large learning rates [19, 20, 21, 22]. Azulay et al. [9] demonstrated that in two-layer linear networks, the relative difference in weight magnitudes between the first and second layer, termed the relative scale in our work, can impact feature learning, with balanced initializations yielding rich learning dynamics, while unbalanced ones tend to induce lazy dynamics. However, as shown in Fig. 1, for nonlinear networks unbalanced initializations can induce both rich and lazy dynamics, creating a complex phase portrait of learning regimes influenced by both overall and relative scale. Building on these observations, our study aims to precisely understand how layer-specific initialization variances and learning rates determine the transition between lazy and rich learning in finite-width networks. Moreover, we endeavor to gain insights into the inductive biases of both regimes, and the transition between them, during training and at interpolation, with the ultimate goal of elucidating how the rich regime acquires features that facilitate generalization.
(a) Overall and relative scale impact feature learning
@@
Figure 1: Unbalanced initializations lead to rapid rich learning and generalization. We follow the experimental setup used in Fig. 1 of Chizat et al. [8] – a wide two-layer student ReLU network $f(x;\theta)=\sum_{i=1}^{h}a_{i}\max(0,w_{i}^{\intercal}x)$ trained on a dataset generated from a narrow two-layer teacher ReLU network. The student parameters are initialized as $w_{i}\sim\text{Unif}(\mathbb{S}^{d-1}(\frac{\tau}{\alpha}))$ and $a_{i}=\pm\alpha\tau$, such that $\tau>0$ controls the overall scale of the function, while $\alpha>0$ controls the relative scale of the first and second layers through the conserved quantity $\delta=\tau^{2}(\alpha^{2}-\alpha^{-2})$. (a) Shows the training trajectories of $|a_{i}|w_{i}$ (color denotes $\mathrm{sgn}(a_{i})$) when $d=2$ for four different settings of $\tau,\delta$. The left plot confirms that small overall scale leads to rich and large overall scale to lazy. The right plot shows that even at small overall scale, the relative scale can move the network between rich and lazy as well. Here an upstream initialization $\delta>0$ shows striking alignment to the teacher (dotted lines), while a downstream initialization $\delta<0$ shows no alignment. (b) Shows the test loss and kernel distance from initialization computed through training over a sweep of $\tau$ and $\delta$ when $d=100$. Lazy learning happens when $\tau$ is large, rich learning happens when $\tau$ is small, and rapid rich learning happens when both $\tau$ is small and $\delta$ is large – an upstream initialization. This initialization also leads to the smallest test loss. See Fig. 10 in Section D.1 for supporting figures.
@@
Our contributions. Our work begins with an exploration of the two-layer single-neuron linear network proposed by Azulay et al. [9] as a minimal model displaying both lazy and rich learning. In Section 3, we derive exact solutions for the gradient flow dynamics with layer-specific learning rates of this model by employing a combination of hyperbolic and spherical coordinate transformations. Alongside recent work by Xu and Ziyin [23] [^1] Xu and Ziyin [23] presented exact NTK dynamics for a linear model trained with one-dimensional data., our analysis stands out as one of the few analytically tractable models for the transition between lazy and rich learning in a finite-width network, marking a notable contribution to the field. Our analysis reveals that the layer-specific initialization variances and learning rates conspire to influence the learning regime through a simple set of conserved quantities that constrain the geometry of learning trajectories. Additionally, it reveals that a crucial aspect of the relative scale overlooked in prior analysis is its directionality. While a balanced initialization results in all layers learning at similar rates, an unbalanced initialization can cause faster learning in either earlier layers, referred to as an upstream initialization, or later layers, referred to as a downstream initialization. Due to the depth-dependent expressivity of layers in a network, upstream and downstream initializations often exhibit fundamentally distinct learning trajectories. In Section 4 we extend our analysis of the relative scale developed in the single-neuron model to more complex linear models with multiple neurons, outputs, and layers and in Section 5 to two-layer nonlinear networks with piecewise linear activation functions. We find that in linear networks, rapid rich learning can only occur from balanced initializations, while in nonlinear networks, upstream initializations can actually accelerate rich learning. Finally, through a series of experiments, we provide evidence that upstream initializations drive feature learning in deep finite-width networks, promote interpretability of early layers in CNNs, reduce the sample complexity of learning hierarchical data, and decrease the time to grokking in modular arithmetic.
@@
Notation. In this work, we consider a feedforward network $f(x;\theta):\mathbb{R}^{d}\to\mathbb{R}^{c}$ parameterized by $\theta\in\mathbb{R}^{m}$. Unless otherwise specified, $c=1$. The network is trained by gradient flow $\dot{\theta}=-{\eta_{\theta}}\cdot\nabla_{\theta}\mathcal{L}(\theta)$, with an initialization $\theta_{0}$ and layer-specific learning rate $\eta_{\theta}\in\mathbb{R}^{m}_{+}$, to minimize the mean squared error $\mathcal{L}(\theta)=\frac{1}{2}\sum_{i=1}^{n}(f(x_{i};\theta)-y_{i})^{2}$ computed over a dataset $\{(x_{1},y_{1}),\dots,(x_{n},y_{n})\}$ of size $n$. We denote the input matrix as $X\in\mathbb{R}^{n\times d}$ with rows $x_{i}\in\mathbb{R}^{d}$ and the label vector as $y\in\mathbb{R}^{n}$. The network’s output $f(x;\theta)$ evolves according to the differential equation, $\partial_{t}f(x;\theta)=\sum_{i=1}^{n}\Theta(x,x_{i};\theta)(y_{i}-f(x_{i};\theta))$, where $\Theta(x,x^{\prime};\theta):\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$ is the Neural Tangent Kernel (NTK), defined as $\Theta(x,x^{\prime};\theta)=\sum_{p=1}^{m}{\eta_{\theta}}_{p}\partial_{\theta_{p}}f(x;\theta)\partial_{\theta_{p}}f(x^{\prime};\theta)$. The NTK quantifies how one gradient step with data point $x^{\prime}$ affects the evolution of the networks’s output evaluated at another data point $x$. When ${\eta_{\theta}}_{p}$ is shared by all parameters, the NTK is the kernel associated with the feature map $\nabla_{\theta}f(x;\theta)\in\mathbb{R}^{m}$. We also define the NTK matrix $K\in\mathbb{R}^{n\times n}$, which is computed across the training data such that $K_{ij}=\Theta(x_{i},x_{j};\theta)$. The NTK matrix evolves from its initialization $K_{0}$ to convergence $K_{\infty}$ through training. Lazy and rich learning exist on a spectrum, with the extent of this evolution serving as the distinguishing factor. Various studies have proposed different metrics to track the evolution of the NTK matrix [24, 25, 26]. We use kernel distance [27], defined as $S(t_{1},t_{2})=1-\langle K_{t_{1}},K_{t_{2}}\rangle/\left(\|K_{t_{1}}\|_{F}\|K_{t_{2}}\|_{F}\right)$, which is a scale invariant measure of similarity between the NTK at two times. In the lazy regime $S(0,t)\approx 0$, while in the rich regime $0\ll S(0,t)\leq 1$.
@@ 2 Related Work
Infinite-width networks. Many recent advancements in understanding the rich regime have come from studying how the initialization variance and layer-wise learning rates should scale in the infinite-width limit to ensure constant movement in the activations, gradients, and outputs. In this limit, analyzing dynamics becomes simpler in several respects: random variables concentrate and quantities will either vanish to zero, remain constant, or diverge to infinity [17]. A set of works used tools from statistical mechanics to provide analytic solutions for the rich population dynamics of two-layer nonlinear neural networks initialized according to the mean field parameterization [56, 57, 58, 59]. These ideas were extended to deeper networks through a tensor program framework, leading to the derivation of maximal update parametrization ($\mu\mathrm{P}$) [16, 18]. The $\mu\mathrm{P}$ parameterization has also been derived through a self-consistent dynamical mean field theory [60] and a spectral scaling analysis [61]. In this study, we focus on finite-width neural networks, but discuss the connection between our work and these width-dependent parameterizations in Section 5.
@@ 3 A Minimal Model of Lazy and Rich Learning with Exact Solutions
Deriving exact solutions in parameter space. We initially assume [^3] We relax this assumption when considering the dynamics of $\beta$ in function space and their implicit bias. whitened input $X^{\intercal}X=\mathbf{I}_{d}$ such that the ordinary least squares solution is $\beta_{*}=X^{\intercal}y$, and the gradient flow dynamics simplify to $\dot{a}=\eta_{a}\left(w^{\intercal}\beta_{*}-a\|w\|^{2}\right)$, $\dot{w}=\eta_{w}\left(a\beta_{*}-a^{2}w\right)$. Notice that $w(t)\in\mathrm{span}(\{w_{0},\beta_{*}\})$, and through training, $w$ aligns in direction to ±$\pm\beta_{*}$ depending on the basin of attraction [^4] The basin is given by $\mathrm{sgn}(a_{0})$ for $\delta\geq 0$ or $\mathrm{sgn}(w_{0}^{\intercal}\beta_{*}+\frac{a_{0}}{2}(\delta+\sqrt{\delta^{2}+4\|\beta_{*}\|^{2}}))$ for $\delta<0$. See A.2.5. the parameters are initialized in. Therefore, we can monitor the dynamics by tracking the hyperbolic geometry between $a$ and $\|w(t)\|$ and the spherical angle between $w(t)$ and $\beta_{*}$. We study the variables $\mu=a\|w\|$, an invariant under the rescale symmetry, and $\phi=\frac{w^{\intercal}\beta_{*}}{\|w\|\|\beta_{*}\|}$, the cosine of the spherical angle. From these two scalar quantities $\mu(t),\phi(t)$ and the initialization $a_{0},w_{0}$, we can determine the trajectory $a(t)$ and $w(t)$ in parameter space. The dynamics for $\mu,\phi$ are given by the coupled nonlinear ODEs,
@@
Amazingly, this system can be solved exactly, as discussed in Section A.2, and shown in Fig. 3. Without delving into the specifics, we can develop an intuitive understanding of the solutions by examining the influence of the relative scale $\delta$.
@@
Identifying regimes of learning in function space. Here we take an alternative route towards understanding the influence of the relative scale by directly examining the dynamics in function space, an analysis strategy we will generalize to broader setups in Sections 4 and 5. The network’s function is determined by the product $\beta=aw$ and governed by the ODE,
@@
where $\rho=X\beta-y$ is the residual. These dynamics can be interpreted as preconditioned gradient flow on the loss in function space where the preconditioning matrix $M$ depends on time through its dependence on $a^{2}$ and $ww^{\intercal}$. Whenever $\|\beta\|\neq 0$, we can express $M$ directly in terms of $\beta$ and $\delta$ as
$M=\frac{\kappa+\delta}{2}\mathbf{I}_{d}+\frac{\kappa-\delta}{2}\frac{\beta\beta^{\intercal}}{\|\beta\|^{2}},$ (4)
where $\kappa=\sqrt{\delta^{2}+4\eta_{a}\eta_{w}\|\beta\|^{2}}$ (see Section A.3 for a derivation). This establishes a self-consistent equation for the dynamics of $\beta$ regulated by $\delta$. Additionally, notice that $M$ characterizes the NTK matrix Eq. 1. Thus, understanding the evolution of $M$ along the trajectory $\beta_{0}$ to $\beta_{*}$ offers a method to discern between lazy and rich learning. Upstream. When $\delta\gg 0$, $M\approx\delta\mathbf{I}_{d}$, and the dynamics of $\beta$ converge to the trajectory of linear regression trained by gradient flow. Along this trajectory the NTK matrix remains constant, confirming the dynamics are lazy. Balanced. When $\delta=0$, $M=\sqrt{\eta_{a}\eta_{w}}\|\beta\|(\mathbf{I}_{d}+\tfrac{\beta\beta^{\intercal}}{\|\beta\|^{2}})$. Here the dynamics balance between following the lazy trajectory and attempting to fit the task by only changing in norm. As a result the NTK changes in both magnitude and direction through training, confirming the dynamics are rich. Downstream. When $\delta\ll 0$, $M\approx|\delta|\tfrac{\beta\beta^{\intercal}}{\|\beta\|^{2}}$, and $\beta$ follows a projected gradient descent trajectory, attempting to reach $\beta_{*}$ in the direction of $\beta_{0}$. Along this trajectory the NTK matrix doesn’t evolve. However, if $\beta_{0}$ is not aligned to $\beta_{*}$, then at some point the dynamics of $\beta$ will slowly align. In this second alignment phase the NTK matrix will change, confirming the dynamics are initially lazy followed by a delayed rich phase. See Section A.3.1 for a derivation of the NTK dynamics $\dot{K}$.